OpenStudy (anonymous):

Jenny left the hardware store at the same as Julio. They drove in opposite directions. Julio drove at a speed of 65mph. After five hours they were 525mi. apart. How fast did Jenny drive?

OpenStudy (blurbendy):

Julio drove a total of 325 miles in 5 hours. Jenny needs to drive 200 miles in order for them to be 525 miles apart. In fives hours, to drive 200 miles IN an hour, Jenny must have been driving at 40 miles per hour.

OpenStudy (anonymous):

okay can you explain or draw out how? I don't get the problem in general

OpenStudy (blurbendy):

Okay, Julio drove 325 miles in 5 hours, correct?

OpenStudy (blurbendy):

The only way he could have done that is if he was driving 65 miles in an hour (65 + 65 + 65 +65 + 65) = 65(5) = 325

OpenStudy (blurbendy):

that's what it means to drive 65 Miles Per Hour

OpenStudy (blurbendy):

However, they were 525 miles apart, 525 - 325 = 200, so Julia needs to drive 200 miles so that they can 525 miles apart.

OpenStudy (blurbendy):

The only way to drive 200 miles in 5 hours is to drive 40 miles PER HOUR (40 + 40 + 40 + 40 + 40) = 40(5) = 200 Understand?

OpenStudy (anonymous):

yes thank you